OpenStudy (anonymous):

he table below shows the distance d(t) in meters that an object travels in t seconds: t (seconds) d(t) (meters) 3 126 4 224 5 350 6 504 What is the average rate of change of d(t) between 3 seconds and 5 seconds, and what does it represent?

OpenStudy (anonymous):

the table below shows the distance d(t) in meters that an object travels in t seconds: t (seconds) d(t) (meters) 3 126 4 224 5 350 6 504 What is the average rate of change of d(t) between 3 seconds and 5 seconds, and what does it represent? 59.5 m/s; it represents the average speed of the object between 3 seconds and 5 seconds 112 m/s; it represents the average speed of the object between 3 seconds and 5 seconds 112 m/s; it represents the average distance traveled by the object between 3 seconds and 5 seconds 59.5 m/s; it represents the average distance traveled by the object between 3 seconds and 5 seconds

OpenStudy (trojanpoem):

\[\frac{ \Delta d }{ \Delta t } = \frac{ 350 - 126 }{ 5-3 } = \] The average change represents the constant average linear speed at this interval.

OpenStudy (anonymous):

224/8

OpenStudy (anonymous):

@TrojanPoem

OpenStudy (trojanpoem):

112

OpenStudy (trojanpoem):

5 - 3 is 2 not 8

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so a or b

OpenStudy (anonymous):

i mean b or c

OpenStudy (anonymous):

i think its b am i right @TrojanPoem

OpenStudy (trojanpoem):

Yeah.

OpenStudy (anonymous):

Thanks